OpenStudy (lxelle):

Differentiate y = x^1/2 ln x

OpenStudy (irishboy123):

take ln for both sides and do it implicitly.

OpenStudy (lxelle):

dy/dx = x^1/2 (1/x) + 1/2x^-1/2 (lnx) right?

OpenStudy (lxelle):

omg i cant really read what you wrote. :P Sorry.

OpenStudy (irishboy123):

agreed. just simplify it.

OpenStudy (lxelle):

So, is mine right?

ganeshie8 (ganeshie8):

thats right!\[\dfrac{dy}{dx} = x^{1/2} (1/x) + 1/2x^{-1/2} (\ln x) \] set that equal to \(0\) and solve \(x\)

OpenStudy (lxelle):

How do i simplify it. D:

OpenStudy (anonymous):

not to butt in, but is this \[y=\sqrt x\ln(x)\]?

OpenStudy (lxelle):

no prob. yes it is

OpenStudy (anonymous):

then i would say rational exponents are not your friends here go with \[\frac{\ln(x)}{2\sqrt{x}}+\frac{\sqrt{x}}{x}\] which is easy to add

OpenStudy (anonymous):

especially since \(\frac{\sqrt{x}}{x}=\frac{1}{\sqrt{x}}\)

OpenStudy (lxelle):

how do i simplify it ffrom there

OpenStudy (lxelle):

@ganeshie8

ganeshie8 (ganeshie8):

set that equal to 0 and solve x

OpenStudy (lxelle):

can't solve it D:

ganeshie8 (ganeshie8):

show ur attempt

OpenStudy (lxelle):

-sqrt x /x = 1/2 x^-1/2 lnx dont know how to continue anymore.

OpenStudy (lxelle):

esp on the right side, how should i simplify it?

ganeshie8 (ganeshie8):

you're almost there ! simply cancel out x^(-1/2) both sides

ganeshie8 (ganeshie8):

Notice : \[\sqrt{x}/x = x^{1/2-1} = x^{-1/2}\]

OpenStudy (lxelle):

but when you move it to the other side would be a negative alrd. D:

ganeshie8 (ganeshie8):

yes

OpenStudy (lxelle):

Hey i got it alrd!!! thanks! :))

ganeshie8 (ganeshie8):

dy/dx = x^1/2 (1/x) + 1/2x^-1/2 (lnx) = 0 x^(-1/2) + 1/2 x^(-1/2) lnx = 0 x^(-1/2) [2 + lnx] = 0 2 + lnx = 0 lnx = -2 x = e^(-2) = 1/e^2

OpenStudy (lxelle):

exsctly.

ganeshie8 (ganeshie8):

good good